OpenStudy (anonymous):

If a satellite orbits the Earth with a period of 120 minutes, determine its altitude above the Earth’s surface. Show all work and answer to three significant digits.

OpenStudy (anonymous):

You can calculate the speed of a geosynchronous orbit of a satellite with the following formula: \[V = \sqrt{\frac{ Gm_E}{ r }}\] and the orbital speed knowing the circumference of a sphere is 2(pi)R and T as the period so the orbital speed is given by: \[\frac{ 2\pi r }{ T }\] Setting them equal to each other and solving for T, period, you know have: \[T = 2 \pi \sqrt{\frac{ r^3 }{ Gm_E }} \] That formula should be familiar now and you can solve for r the radius of earth+ atmosphere/altitude of satellite. You will need radius of the Earth to subtract it from the answer for r that you find. I believe the Earth's radius is:\[6.38 \times 10^6 m\] Keep watch for units.

OpenStudy (anonymous):

i manipulated to get r=(T^2GmE/2π)^1/3 and plugged in the values =[(7200s)^2(6.67x10^-11)/2π]^1/3 an i got 0.0819 which seems wrong to me @johnbc

OpenStudy (anonymous):

You forgot to also square the 2 pi.

OpenStudy (anonymous):

@lolo123

OpenStudy (anonymous):

thank u so much!

OpenStudy (anonymous):

My pleasure